Solve the equation by first clearing the fraction or decimals.

52. 11.6b - 5 = 9.5
Any help with this equation would be Greatly Aprreciated.

Answer by checkley71(8403)

11.6B-5=9.5
116B-50=95
116B=95+50
116B=145
B=145/116
B=1.25 ANSWER.